OpenStudy (inowalst):

What is the simplified form of \[\frac{ x+4 }{ x^2+x-12 }\times \frac{ x+4 }{ x^2+8x+16 }\] ?

OpenStudy (zale101):

Can you factor the denominators? If you factored them, you can get some cancelings..

OpenStudy (inowalst):

That's what I need help with. Factoring them out.

OpenStudy (inowalst):

\[\frac{ 1 }{ (x+3)(x-4) }\] \[\frac{ 1 }{ (x-4)^2 }\] \[\frac{ 1 }{ (x+4)^2 }\] \[\frac{ 1 }{ (x-3)(x+4) }\]

Nnesha (nnesha):

that is for left side bottom one ?? which is x^2 + x -12

OpenStudy (inowalst):

Yes.

Nnesha (nnesha):

okay so for that easy method is to find two number if you add them you should get middle number which is 1 and if you multiply that two numbers you should get -12 so -4 times 3 = ?? -4+ 3 = ?

OpenStudy (inowalst):

-12 -1

Nnesha (nnesha):

@Zale101 look it there she is back!!

Nnesha (nnesha):

yes -4 +3 = -1 but you should have positive one so play with number and see how you get positive one

OpenStudy (inowalst):

With any number?

Nnesha (nnesha):

you can factor -12

Nnesha (nnesha):

you have -4 and 3 which is almost right play with signs and see if you get positive one

OpenStudy (inowalst):

4+(-3)=1

Nnesha (nnesha):

yes that's right because your leading coefficient is one so you can just right it as (x +4) and (x-3)

OpenStudy (inowalst):

Ohh okay.

Nnesha (nnesha):

now factor right bottom side again because your leading coefficient is 1 so you can just find two number if you multiply them you should get positive 16 and if you add or subtract them you should get positive 8

OpenStudy (inowalst):

Two same numbers to get 16 or any?

Nnesha (nnesha):

some times it can be same but not always

OpenStudy (inowalst):

Okay.

Nnesha (nnesha):

okay lets do it this way  factor 16 and factor x^2